OpenStudy (anonymous):

Demand Equation The price (in dollars) and the quantity x sold of a certain product obey the demand equation p= -1/10x+150, 0≤x≤1500 Revenue is x*p. (a) Express the revenue R as a function of x. (b) What quantity x maximizes revenue? What is the maximum revenue?

OpenStudy (mathstudent55):

\(p = -\dfrac{1}{10}x + 150 \) \(R(x) = px = (-\dfrac{1}{10}x + 150 )x \) \( R(x) =-\dfrac{1}{10}x^2 + 150x\) This is a parabola that opens downward. To find the value of x that maximizes R, find the equation of the axis of symmetry. Then plug in that value of x into the equation for R.

OpenStudy (anonymous):

thank you

OpenStudy (mathstudent55):

wlcm